OpenStudy (anonymous):

what is the simplified form of (0.00038 x 10^3)^2 in standard notation?

OpenStudy (anonymous):

expand the bracket first..you will get \[0.0001444 \times 10^{6}\] then simplify until you get \[1.444 \times 10^{-4} \times 10^{6}\] and the answer is \[1.444 \times 10^{2}\]

OpenStudy (anonymous):

thanks!